OpenStudy (anonymous):

A car is traveling at a constant velocity. After two hours the car has traveled 90 miles. After five hours the car has traveled 225 miles. Assuming a linear equation fits this situation, how far will the car have traveled after seven hours?

OpenStudy (texaschic101):

2/90 = 1/x (2 hrs car traveled 90 miles = 1 hour traveled x miles) cross multiply 2x = 90 x = 45 So the car travels 45 miles per hour. distance traveled = 45 times x (x being the number of miles) D = (45)(x) D = (45)(7) D = 315

OpenStudy (anonymous):

so the answer is 315

OpenStudy (texaschic101):

yes.....the linear equation is : D = 45x (where D is distance and x is number of miles. 315 is the answer to the equation. I think you will need both the equation and the answer.